OpenStudy (anonymous):

What is the slope of the line given by the equation 2x + 3y = 6?

OpenStudy (anonymous):

you have a choice you can solve for \(y\) via \[2x+3y=6\\ 3y=-2x+6\\ y=-\frac{2}{3}x+2\] now it looks like \(y=mx+b\) with slope \(m\) i.e. the slope is \(-\frac{2}{3}\)

OpenStudy (anonymous):

or you can remember that the slope of \[ax+by=c\] is always \[m=-\frac{a}{b}\]